# Federated Hierarchical Tensor Networks: a Collaborative Learning Quantum AI-Driven Framework for Healthcare

### v1

@@ I INTRODUCTION
In parallel, tensor networks (TNs) are considered promising candidates for quantum machine learning (QML) architectures [18]. TNs provide a compact and efficient representation of quantum states, using tensor structures that encapsulate substantial amounts of correlated information. TNs enable the application of individual local operations on each tensor node, eliminating the need to compute the entire tensor. Tensors can be combined through contraction over connected indices or broken down into multiple interconnected tensors. For a more in-depth technical exploration of TNs, readers can delve into a comprehensive introduction [19] and review articles that focus on specific architectures [20, 21, 22]. Matrix Product State (MPS) is the simplest mathematical framework for describing quantum states in many-body systems [23, 24]. The key idea behind MPS is that the entanglement in a many-body quantum state is often localized, meaning that each tensor is primarily entangled with its neighbors. It is a widely studied one-dimensional TN layout in which the level of entanglement remains constant. In hierarchical tensor networks (HTNs), input or output tensors are pooled across several internal layers rather than directly interconnected. The tree tensor network (TTN), a specific form of HTN, is designed to efficiently represent high-dimensional tensors by decomposing them into a sequence of lower-dimensional tensors interconnected hierarchically. One of the key advantages of TTN is its ability to capture long-range correlations in high-dimensional data while maintaining a low computational cost [18]. Another more complex TN layout is the multiscale entanglement renormalization ansatz (MERA) that goes beyond TTN in capturing symmetry and handling higher entanglement at different length scale, making it a valuable tool for studying complex quantum systems [25, 18], as shown in Fig. 2(a-c).
@@ IV.2 Implementation details
All methods use the Adaptive Moment Estimation (Adam) optimizer; the learning rate of global server and local models is set to 0.001, mini-batch size of 8, and weight decay of 1e-4. All the experiments are performed using a quantum programming framework, TensorFlow Quantum (TFQ), by Google [38]. To implement classical MPS, we used the Tensor Network Library [39]. For differential privacy, we adopted a differentially private DPAdam optimizer, clipping C = 1.0, and noise level $\epsilon=[0.2,0.4,0.6,0.8]$. The hardware and software utilized in the paper are described within section. IV
@@ IV.6 Global model
The global model is responsible for coordinating and aggregating the local models from each hospital. The aggregation process ensures that the knowledge gained by individual hospitals is combined to improve the overall QTN global model. Upon completion of local training for communication round (r), each hospital or client communicates its local model parameters ($\theta_{h}$) to the global server, maintaining the privacy of the local data within the hospitals that generated it. Finally, the aggregation of all local model updates takes place using a federated averaging algorithm as $\theta^{r}_{s}\leftarrow\theta^{r}_{h}-\eta_{s}\dfrac{1}{H}\sum_{i=1}^{H}\theta^{r+1}_{h}$. The goal is to achieve an optimal global FedQTN model by minimizing the global loss function with a server’s learning rate $\eta_{s}$. The global model broadcasts the updated parameters to all participants. This iterative process adjusts the aggregated model through the local training procedure across hospitals until the quantum model reaches a specified learning accuracy.
@@
We implemented the experiments on the Anvil system at Purdue University through allocation CIS240105 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program [40]. Regarding software, the experiments were performed using TensorFlow 2.7.0 in Python 2.8.

### v2

@@ I INTRODUCTION
Recent strides in federated learning empower the construction of complex classical machine learning models trained across decentralized systems. The effective integration of FL in healthcare stands poised to significantly transform precision medicine on a broad scale and can improve clinical workflow, while also upholding stringent standards of privacy and governance [3]. Noteworthy examples illuminate the promising results of federated learning algorithms in healthcare applications including detecting rare cancer disease of glioblastoma across six continents [4], identifying brain anomalies across six institutions [5], integrating differential privacy to clinical and epidemiological research [6], safeguarding privacy in medical imaging [7], conducting a multi-national study to detect COVID-19 lung abnormalities in CT scans [8], utilizing FL deep generative models for healthcare [9], ensuring privacy in medical image analysis [10], diagnosing COVID-19 with collaborative FL [11], and numerous other breakthroughs. FL stands out as the foremost privacy preservation technique of the next generation, gaining substantial traction across industries and notably within medical AI applications.
@@
In parallel, tensor networks (TNs) are considered promising candidates for quantum machine learning (QML) architectures [27]. TNs provide a compact and efficient representation of quantum states, using tensor structures that encapsulate substantial amounts of correlated information. TNs enable the application of individual local operations on each tensor node, eliminating the need to compute the entire tensor. Tensors can be combined through contraction over connected indices or broken down into multiple interconnected tensors. For a more in-depth technical exploration of TNs, readers can delve into a comprehensive introduction [28] and review articles that focus on specific architectures [29, 30, 31]. Matrix Product State (MPS) is the simplest mathematical framework for describing quantum states in many-body systems [32, 33]. The key idea behind MPS is that the entanglement in a many-body quantum state is often localized, meaning that each tensor is primarily entangled with its neighbors. It is a widely studied one-dimensional TN layout in which the level of entanglement remains constant. In hierarchical tensor networks (HTNs), input or output tensors are pooled across several internal layers rather than directly interconnected. The tree tensor network (TTN), a specific form of HTN, is designed to efficiently represent high-dimensional tensors by decomposing them into a sequence of lower-dimensional tensors interconnected hierarchically. One of the key advantages of TTN is its ability to capture long-range correlations in high-dimensional data while maintaining a low computational cost [27]. Another more complex TN layout is the multiscale entanglement renormalization ansatz (MERA) that goes beyond TTN in capturing symmetry and handling higher entanglement at different length scales, making it a valuable tool for studying complex quantum systems [34, 27], as shown in Fig. 2(a-c).
@@ IV.2 Implementation details
All methods use the Adaptive Moment Estimation (Adam) optimizer; the learning rate of global server and local models is set to 0.001, mini-batch size of 8, and weight decay of 1e-4. All the experiments are performed using a quantum programming framework, TensorFlow Quantum (TFQ), by Google [47]. To implement classical MPS, we used the Tensor Network Library [48]. For differential privacy, we adopted a differentially private DPAdam optimizer, clipping C = 1.0, and noise level $\epsilon=[0.2,0.4,0.6,0.8]$. The hardware and software utilized in the paper are described within the section. IV
@@ IV.6 Global model
The global model is responsible for coordinating and aggregating the local models from each hospital. The aggregation process ensures that the knowledge gained by individual hospitals is combined to improve the overall QTN global model. Upon completion of local training for communication round (r), each hospital or client communicates its local model parameters ($\theta_{h}$) to the global server, maintaining the privacy of the local data within the hospitals that generated it. Finally, the aggregation of all local model updates takes place using a federated averaging algorithm as
$\theta^{r}_{s}\leftarrow\theta^{r}_{h}-\eta_{s}\dfrac{1}{H}\sum_{i=1}^{H}\theta^{r+1}_{h}$ (5)
. The goal is to achieve an optimal global FedQTN model by minimizing the global loss function with a server’s learning rate $\eta_{s}$. The global model broadcasts the updated parameters to all participants. This iterative process adjusts the aggregated model through the local training procedure across hospitals until the quantum model reaches a specified learning accuracy.
@@
We implemented the experiments on the Anvil system at Purdue University through allocation CIS240105 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program [49]. Regarding software, the experiments were performed using TensorFlow 2.7.0 [50] in Python 2.8.